OpenStudy (anonymous):

A jar holds 15 red pencils and 10 blue pencils. What is the probability of drawing two red pencils from the jar

OpenStudy (anonymous):

there are two different answers depending on whether or not you are doing this with replacement (you take one out and put it back in), or without replacement (you take one out, and it is no longer in the jar), does your question specify which replacement is used?

OpenStudy (dumbcow):

i think without replacement is implied

OpenStudy (anonymous):

No replacement

OpenStudy (anonymous):

then the probability is 15/25 x 14/24

OpenStudy (anonymous):

then what you get without replacement is as follows, you want to draw a red pencil out on the first try, so\[\frac{ number of red pencils }{ total pencils }=\frac{15}{25}\]on the second draw, you have to remember that you just pulled out a red pencil, so\[\frac{ 14 }{ 24 }\]then you just multiply these together to get your probability

OpenStudy (anonymous):

Thank you you've been a great help